# Nonparametric Proportional Hazards Model with Differential Regularization Applied to Spatial Survival Data

### v4

@@ Appendix C Proof of Theorem 2
which establishes Condition 3.5(i).
Proof of condition 3.5(ii)
Let $K(n)=\dim(\mathcal{H}_{n})$ and $K_{\Theta}(n)=b+K(n)$. Empirical-process entropy bounds for Lipschitz parameterizations yield
@@
Condition 3.6 The sample $\{Z_{t}\}_{t=1}^{n}=\{(Y_{t},\delta_{t},\bm{X}_{t},\bm{P}_{t})\}_{t=1}^{n}$ is i.i.d. or $m$-dependent.
Proof of Condition 3.6
By assumption the observations are i.i.d.
Condition 3.7 There exists a constant $C_{1}>0$ (independent of $n$) such that for all sufficiently small $\varepsilon>0$ and for every sieve $\Theta_{n}$,
@@
Proof of Condition 3.7
Fix $\theta=(\bm{\beta},h)$ and write $\eta=(\bm{u},g)=(\bm{\beta}-\bm{\beta}_{0},h-h_{0})$. Decompose
@@
uniformly over $d(\theta,\theta_{0})\leq\varepsilon$ and $\theta\in\Theta_{n}$.
Condition 3.8 For any $\delta>0$ there exists $s\in(0,2)$ and a random variable $U(Z)$ with $\mathbb{E}\big[(U(Z))^{\gamma}\big]\leq C_{2}$ for some $\gamma\geq 2$ such that
@@
Proof of Condition 3.8
The Lipschitz estimate obtained in Condition 3.7 yields, for all $\theta$ in a $\delta$–ball around $\theta_{0}$,
@@
Thus the inequality holds with $s=1$ and $U(Z)\equiv C^{\prime}$. Since $C^{\prime}$ is deterministic, $\mathbb{E}\big[(U(Z))^{\gamma}\big]=C^{\prime\gamma}<\infty$ for any $\gamma\geq 2$, which verifies Condition 3.8. ∎
@@ Appendix D Proof of Theorem 3
We apply Theorem 4.3 of chen2007large verifying Conditions 4.1–4.5 for the functional $f(\theta)=\bm{a}^{\top}\bm{\beta}$ and then use Cramér–Wold to obtain the joint limit for $\bm{\beta}$; here $\bm{a}\in\mathbb{R}^{b}$ is arbitrary.
By Theorem 4.3 of chen2007large and our rate assumption $\|\hat{\theta}_{n}-\theta_{0}\|^{2}=o_{p}(n^{-1/2})$ (implied by $\varepsilon_{n}^{2}=o(n^{-1/2})$ and $\|\hat{\theta}_{n}-\theta_{0}\|=O_{p}(\varepsilon_{n})$), it holds
@@
We now prove Condition 4.1-4.5 of chen2007large. Write $\theta=(\bm{\beta},h)$, $\theta_{0}=(\bm{\beta}_{0},h_{0})$, and define the population unpenalized partial log-likelihood contribution
@@
Condition 4.1 $(i)$ There is $\omega>0$ such that
$|f(\theta)-f(\theta_{o})-\frac{\partial f(\theta_{o})}{\partial\theta}[\theta-\theta_{o}]|=O(\|\theta-\theta_{o}\|^{\omega})$
uniformly in $\theta\in\Theta_{n}$ with $\|\theta-\theta_{o}\|=o(1)$; $(ii)$ $\left\|\frac{\partial f(\theta_{o})}{\partial\theta}\right\|<\infty$; $(iii)$ there is $\pi_{n}v^{*}\in\Theta_{n}$ such that
$\|\pi_{n}v^{*}-v^{*}\|\times\|\hat{\theta}_{n}-\theta_{o}\|=o_{P}(n^{-1/2}).$
Proof of Condition 4.1
Condition 4.1 $(i)$ and $(ii)$ hold with $\omega=2$ because $f(\theta)=\bm{a}^{\top}\bm{\beta}$ is linear and $\big\|\partial f(\theta_{0})/\partial\theta\big\|<\infty$. For 4.1 $(iii)$, by the FEM approximation bound (4), $\|\pi_{n}v_{\bm{a}}^{*}-v_{\bm{a}}^{*}\|=O(\eta(n))$, and $\|\hat{\theta}_{n}-\theta_{0}\|=O_{p}(\varepsilon_{n})$ by Theorem 2. Hence
@@
verifying 4.1 $(iii)$.
Condition 4.2 Let $\mu_{n}(g(Z))=n^{-1}\sum_{i=1}^{n}g(z_{i})-E[g(Z_{i})]$ denote the empirical process indexed by the function a function $g$. Then,
@@
Proof of Condition 4.2
The map $\theta\mapsto\ell(\theta,Z)$ is twice Fréchet differentiable in a neighborhood of $\theta_{0}$ (Lemma 4); thus pathwise differentiability holds. Using the Donsker property and (19)–(20), we obtain
@@
which is Condition 4.2 and hence implies 4.2; see chen2007large.
Condition 4.3
$K(\theta_{o},\hat{\theta}_{n})-K(\theta_{o},\hat{\theta}_{n}\pm\varepsilon_{n}\pi_{n}v^{*})=\pm\varepsilon_{n}(\hat{\theta}_{n}-\theta_{o},\pi_{n}v^{*})+o(n^{-1}),$
where $K(\theta_{0},\theta)\equiv\mathbb{E}[l(\theta_{0},Z_{i})-l(\theta,Z_{i})]$.
Proof of Condition 4.3
Taylor-expanding the population criterion $Q(\theta)=P\ell(\theta,\cdot)$ along the line $\theta_{0}\pm\varepsilon_{n}\pi_{n}v_{\bm{a}}^{*}$ and using that $D^{2}Q(\theta_{0})[\xi_{1},\xi_{2}]=\langle\xi_{1},\xi_{2}\rangle$ with linear remainder $o(\varepsilon_{n}^{2})$, we get
@@
which is Condition 4.3; hence 4.3 holds. The penalty contributes nothing to these equalities in the $\bm{\beta}$ direction, and its contribution along $g$ is of order $o(\lambda_{n}\varepsilon_{n})=o(n^{-1/2})$.
Condition 4.4 $(i)$ $\mu_{n}\left(\frac{\partial l(\theta_{o},Z)}{\partial\theta}[\pi_{n}v^{*}-v^{*}]\right)=o_{P}(n^{-1/2});$ $(ii)$ $E\left\{\frac{\partial l(\theta_{o},Z)}{\partial\theta}[\pi_{n}v^{*}]\right\}=o(n^{-1/2}).$
Proof of Condition 4.4
Because $\|\pi_{n}v_{\bm{a}}^{*}-v_{\bm{a}}^{*}\|=O(\eta(n))$ and the score map $\xi\mapsto\dot{\ell}(\theta_{0},Z)[\xi]$ is square-integrable and continuous in $\xi$ under $\langle\cdot,\cdot\rangle$, we have
@@
since $\mathbb{E}[\dot{\ell}(\theta_{0},Z)[v_{\bm{a}}^{*}]]=0$ by the definition of the representer and the score identity.
Condition 4.5
$n^{1/2}\,\mu_{n}\left(\frac{\partial l(\theta_{o},Z)}{\partial\theta}[v^{*}]\right)\overset{d}{\to}\mathcal{N}(0,\sigma_{v^{*}}^{2}),\quad\text{with }\sigma_{v^{*}}^{2}>0.$
Proof of Condition 4.5
The i.i.d. assumption and finite variance imply the CLT for the score in direction $v_{\bm{a}}^{*}$:
@@
which is Condition 4.5. Note that the penalty does not enter the score, and the $S_{n}^{(0)}$ instead of $s^{(0)}$ replacement alters the criterion by $O_{p}(\delta_{n})$ uniformly by (20); combined with the fact that $\delta_{n}\epsilon_{n}=o(n^{-1/2})$, its contribution to the linearization is negligible for the $\sqrt{n}$-asymptotics. ∎
@@ Appendix E Lemmas
$C_{*}:=M_{\mathcal{B}}M_{\mathcal{X}}+M_{\infty}.$
@@
with $L_{*}:=(1+e^{2C_{*}})\max\{M_{\mathcal{X}},1\}$, which is deterministic and finite. Taking expectations and using $\delta\leq 1$,
@@
$s^{(k)}(\theta,y)[\eta_{1},\ldots,\eta_{k}]:=\mathbb{E}\!\Bigl[\mathbf{1}\{Y^{\prime}\!\geq y\}\,\exp\bigl(\bm{X}^{\prime\!\top}\bm{\beta}+h(\bm{P}^{\prime})\bigr)\,\prod_{r=1}^{k}\bigl(\bm{X}^{\prime\!\top}\bm{u}_{r}+g_{r}(\bm{P}^{\prime})\bigr)\Bigr],$
@@
Proof.
Assumption 2(i) gives $\|\bm{X}\|\leq M_{\mathcal{X}}$ a.s., $\mathcal{B}$ is compact so $\|\bm{\beta}\|\leq M_{\mathcal{B}}$, and the Sobolev–Morrey embedding yields a uniform bound $\|h\|_{\infty}\leq C_{H}M_{\mathcal{H}}$ for all $h\in\mathcal{H}$. Let $C_{*}:=M_{\mathcal{B}}M_{\mathcal{X}}+C_{H}M_{\mathcal{H}}$. Then for any $\theta=(\bm{\beta},h)$ and $y\geq 0$,
@@
Moreover, for $\eta=(\bm{u},g)$ write $L(\eta):=M_{\mathcal{X}}\|\bm{u}\|+\|g\|_{\infty}$. By the mean–value theorem and boundedness of the exponential,
@@
Again, the integrand is dominated by a deterministic constant times $L(\eta_{1})L(\eta_{2})$, so the limit passes through the expectation. Bilinearity and continuity in $(\eta_{1},\eta_{2})$ are immediate from the bounds, which also give the $o(\|\eta_{2}\|)$ remainder uniformly in $\eta_{1}$. Thus $Q$ is $C^{2}$ Fréchet on the interior of $\Theta$ with the stated $DQ$ and $D^{2}Q$. ∎

### v5

@@ Appendix C Proof of Theorem 2
which establishes Condition 3.5(i). $\square$
Proof of condition 3.5(ii): Let $K(n)=\dim(\mathcal{H}_{n})$ and $K_{\Theta}(n)=b+K(n)$. Empirical-process entropy bounds for Lipschitz parameterizations yield
@@
$\square$
Condition 3.6. The sample $\{Z_{t}\}_{t=1}^{n}=\{(Y_{t},\delta_{t},\bm{X}_{t},\bm{P}_{t})\}_{t=1}^{n}$ is i.i.d. or $m$-dependent.
Proof of Condition 3.6 By assumption the observations are i.i.d. $\square$
Condition 3.7. There exists a constant $C_{1}>0$ (independent of $n$) such that for all sufficiently small $\varepsilon>0$ and for every sieve $\Theta_{n}$,
@@
Proof of Condition 3.7: Fix $\theta=(\bm{\beta},h)$ and write $\eta=(\bm{u},g)=(\bm{\beta}-\bm{\beta}_{0},h-h_{0})$. Decompose
@@
uniformly over $d(\theta,\theta_{0})\leq\varepsilon$ and $\theta\in\Theta_{n}$. $\square$
Condition 3.8. For any $\delta>0$ there exists $s\in(0,2)$ and a random variable $U(Z)$ with $\mathbb{E}\big[(U(Z))^{\gamma}\big]\leq C_{2}$ for some $\gamma\geq 2$ such that
@@
Proof of Condition 3.8 The Lipschitz estimate obtained in Condition 3.7 yields, for all $\theta$ in a $\delta$–ball around $\theta_{0}$,
@@
Thus the inequality holds with $s=1$ and $U(Z)\equiv C^{\prime}$. Since $C^{\prime}$ is deterministic, $\mathbb{E}\big[(U(Z))^{\gamma}\big]=C^{\prime\gamma}<\infty$ for any $\gamma\geq 2$, which verifies Condition 3.8. $\square$
@@ Appendix D Proof of Theorem 3
We apply Theorem 4.3 of Chen [2007] verifying Conditions 4.1–4.5 for the functional $f(\theta)=\bm{a}^{\top}\bm{\beta}$ and then use Cramér–Wold to obtain the joint limit for $\bm{\beta}$; here $\bm{a}\in\mathbb{R}^{b}$ is arbitrary.
By Theorem 4.3 of Chen [2007] and our rate assumption $\|\hat{\theta}_{n}-\theta_{0}\|^{2}=o_{p}(n^{-1/2})$ (implied by $\varepsilon_{n}^{2}=o(n^{-1/2})$ and $\|\hat{\theta}_{n}-\theta_{0}\|=O_{p}(\varepsilon_{n})$), it holds
@@
We now prove Condition 4.1-4.5 of Chen [2007]. Write $\theta=(\bm{\beta},h)$, $\theta_{0}=(\bm{\beta}_{0},h_{0})$, and define the population unpenalized partial log-likelihood contribution
@@
Condition 4.1. $(i)$ There is $\omega>0$ such that $|f(\theta)-f(\theta_{o})-\frac{\partial f(\theta_{o})}{\partial\theta}[\theta-\theta_{o}]|=O(\|\theta-\theta_{o}\|^{\omega})$ uniformly in $\theta\in\Theta_{n}$ with $\|\theta-\theta_{o}\|=o(1)$; $(ii)$ $\left\|\frac{\partial f(\theta_{o})}{\partial\theta}\right\|<\infty$; $(iii)$ there is $\pi_{n}v^{*}\in\Theta_{n}$ such that $\|\pi_{n}v^{*}-v^{*}\|\times\|\hat{\theta}_{n}-\theta_{o}\|=o_{P}(n^{-1/2}).$
Proof of Condition 4.1: Condition 4.1 $(i)$ and $(ii)$ hold with $\omega=2$ because $f(\theta)=\bm{a}^{\top}\bm{\beta}$ is linear and $\big\|\partial f(\theta_{0})/\partial\theta\big\|<\infty$. For 4.1 $(iii)$, by the FEM approximation bound (4), $\|\pi_{n}v_{\bm{a}}^{*}-v_{\bm{a}}^{*}\|=O(\eta(n))$, and $\|\hat{\theta}_{n}-\theta_{0}\|=O_{p}(\varepsilon_{n})$ by Theorem 2. Hence
@@
verifying 4.1 $(iii)$. $\square$
Condition 4.2. Let $\mu_{n}(g(Z))=n^{-1}\sum_{i=1}^{n}g(z_{i})-E[g(Z_{i})]$ denote the empirical process indexed by the function a function $g$. Then,
@@
Proof of Condition 4.2: The map $\theta\mapsto\ell(\theta,Z)$ is twice Fréchet differentiable in a neighborhood of $\theta_{0}$ (Lemma 4); thus pathwise differentiability holds. Using the Donsker property and (19)–(20), we obtain
@@
which is Condition 4.2 and hence implies 4.2; see Chen [2007]. $\square$
Condition 4.3. $K(\theta_{o},\hat{\theta}_{n})-K(\theta_{o},\hat{\theta}_{n}\pm\varepsilon_{n}\pi_{n}v^{*})=\pm\varepsilon_{n}(\hat{\theta}_{n}-\theta_{o},\pi_{n}v^{*})+o(n^{-1}),$ where $K(\theta_{0},\theta)\equiv\mathbb{E}[l(\theta_{0},Z_{i})-l(\theta,Z_{i})]$.
Proof of Condition 4.3: Taylor-expanding the population criterion $Q(\theta)=P\ell(\theta,\cdot)$ along the line $\theta_{0}\pm\varepsilon_{n}\pi_{n}v_{\bm{a}}^{*}$ and using that $D^{2}Q(\theta_{0})[\xi_{1},\xi_{2}]=\langle\xi_{1},\xi_{2}\rangle$ with linear remainder $o(\varepsilon_{n}^{2})$, we get
@@
which is Condition 4.3; hence 4.3 holds. The penalty contributes nothing to these equalities in the $\bm{\beta}$ direction, and its contribution along $g$ is of order $o(\lambda_{n}\varepsilon_{n})=o(n^{-1/2})$. $\square$
Condition [.4.4] $(i)$ $\mu_{n}\left(\frac{\partial l(\theta_{o},Z)}{\partial\theta}[\pi_{n}v^{*}-v^{*}]\right)=o_{P}(n^{-1/2});$ $(ii)$ $E\left\{\frac{\partial l(\theta_{o},Z)}{\partial\theta}[\pi_{n}v^{*}]\right\}=o(n^{-1/2}).$
Proof of Condition 4.4: Because $\|\pi_{n}v_{\bm{a}}^{*}-v_{\bm{a}}^{*}\|=O(\eta(n))$ and the score map $\xi\mapsto\dot{\ell}(\theta_{0},Z)[\xi]$ is square-integrable and continuous in $\xi$ under $\langle\cdot,\cdot\rangle$, we have
@@
since $\mathbb{E}[\dot{\ell}(\theta_{0},Z)[v_{\bm{a}}^{*}]]=0$ by the definition of the representer and the score identity. $\square$
Condition 4.5. $n^{1/2}\,\mu_{n}\left(\frac{\partial l(\theta_{o},Z)}{\partial\theta}[v^{*}]\right)\overset{d}{\to}\mathcal{N}(0,\sigma_{v^{*}}^{2}),\quad\text{with }\sigma_{v^{*}}^{2}>0.$
Proof of Condition 4.5: The i.i.d. assumption and finite variance imply the CLT for the score in direction $v_{\bm{a}}^{*}$:
@@
which is Condition 4.5. Note that the penalty does not enter the score, and the $S_{n}^{(0)}$ instead of $s^{(0)}$ replacement alters the criterion by $O_{p}(\delta_{n})$ uniformly by (20); combined with the fact that $\delta_{n}\epsilon_{n}=o(n^{-1/2})$, its contribution to the linearization is negligible for the $\sqrt{n}$-asymptotics. $\square$
@@ Appendix E Lemmas
$C_{*}=M_{\mathcal{B}}M_{\mathcal{X}}+M_{\infty}.$
@@
with $L_{*}=(1+e^{2C_{*}})\max\{M_{\mathcal{X}},1\}$, which is deterministic and finite. Taking expectations and using $\delta\leq 1$,
@@
$s^{(k)}(\theta,y)[\eta_{1},\ldots,\eta_{k}]=\mathbb{E}\!\Bigl[\mathbf{1}\{Y^{\prime}\!\geq y\}\,\exp\bigl(\bm{X}^{\prime\!\top}\bm{\beta}+h(\bm{P}^{\prime})\bigr)\,\prod_{r=1}^{k}\bigl(\bm{X}^{\prime\!\top}\bm{u}_{r}+g_{r}(\bm{P}^{\prime})\bigr)\Bigr],$
@@
Assumption 2 (i) gives $\|\bm{X}\|\leq M_{\mathcal{X}}$ a.s., $\mathcal{B}$ is compact so $\|\bm{\beta}\|\leq M_{\mathcal{B}}$, and the Sobolev–Morrey embedding yields a uniform bound $\|h\|_{\infty}\leq C_{H}M_{\mathcal{H}}$ for all $h\in\mathcal{H}$. Let $C_{*}=M_{\mathcal{B}}M_{\mathcal{X}}+C_{H}M_{\mathcal{H}}$. Then for any $\theta=(\bm{\beta},h)$ and $y\geq 0$,
@@
Moreover, for $\eta=(\bm{u},g)$ write $L(\eta)=M_{\mathcal{X}}\|\bm{u}\|+\|g\|_{\infty}$. By the mean–value theorem and boundedness of the exponential,
@@
Again, the integrand is dominated by a deterministic constant times $L(\eta_{1})L(\eta_{2})$, so the limit passes through the expectation. Bilinearity and continuity in $(\eta_{1},\eta_{2})$ are immediate from the bounds, which also give the $o(\|\eta_{2}\|)$ remainder uniformly in $\eta_{1}$. Thus $Q$ is $C^{2}$ Fréchet on the interior of $\Theta$ with the stated $DQ$ and $D^{2}Q$.